OpenStudy (anonymous):

find the common ratio of the sequence 2, -10, 50, -250

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

\[\frac{-10}{2}\] or \[\frac{50}{-10}\] or \[\frac{-250}{50}\]your choice